Break-Even Sales

BeerBev, Inc., reported the following operating information for a recent year:

Net sales $6,528,000
Cost of goods sold $1,632,000
Selling, general and administration 816,000
$2,448,000
Income from operations $ 4,080,000*

*Before special items

In addition, assume that BeerBev sold 51,000 barrels of beer during the year. Assume that variable costs were 75% of the cost of goods sold and 50% of selling, general and administration expenses. Assume that the remaining costs are fixed. For the following year, assume that BeerBev expects pricing, variable costs per barrel, and fixed costs to remain constant, except that new distribution and general office facilities are expected to increase fixed costs by $24,500.

When computing the cost per unit amounts for the break-even formula, round to two decimal places. If required, round your final answer to one decimal place.

a. Compute the break-even number of barrels for the current year.
barrels

b. Compute the anticipated break-even number of barrels for the following year.
barrels
Answer #1

variable cost = (1632000*75%+816000*50%) = 1632000

Contribution margin per barrel = (6528000-1632000)/51000 = 96

Fixed cost = 2448000-1632000 = 816000

a) Break even = Fixed cost/Contribution margin per barrels = 816000/96 = 8500 Barrel

b) Break even = (816000+24500)/96 = 8755.2 barrel